# Graph Neural Network Enhanced Retrieval for Question Answering of LLMs

### v1

@@ 5 Conclusion
In this paper, we proposed GNN-Ret, an effective method to enhance retrieval for QA of LLMs by exploiting the inherent relatedness between passages on a graph of passages. By extending GNN-Ret to multi-hop reasoning questions, we proposed RGNN-Ret, which enhances the retrieval for subquestions through the interconnection between graphs of passages across steps. The experiments clearly demonstrated the superiority of both GNN-Ret and RGNN-Ret over baselines, highlighting the effectiveness of leveraging the relatedness between passages to enhance the retrieval process. From these advantages, we believe that the incorporation of graph-based representations and the exploitation of passage relatedness can open up new avenues of research in the field of LLMs in understanding structural documents and answering complex questions.
@@ A.1 Training of GNN
$\bar{d}_{Y}^{L}=\frac{1}{|\mathcal{S}_{Y}|}\sum_{j\in\mathcal{S}_{Y}}h_{j}^{L},\quad\bar{d}_{O/Y}^{L}=\frac{1}{|\mathcal{S}_{O/Y}|}\sum_{o\in\mathcal{S}_{O/Y}}h_{o}^{L},$ (8)
@@
We update the parameters of GNN using gradient descent, and the gradient of $\ell$ with respect to $\alpha^{l}$ is given by:
$\frac{\partial\ell}{\partial\alpha^{l}}=\left\{\begin{aligned} 0,\quad\ell<=0\\ f_{\alpha^{l}}(L,\mathcal{S}_{Y}),\quad\ell>0\end{aligned}\right.,$ (10)
where
$f_{\alpha^{l}}(L,\mathcal{S}_{Y})$ $=\frac{\sum_{j\in\mathcal{S}_{Y}}(h_{j}^{l-1}-m_{j}^{l})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{j}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}_{Y}|}$ (11)
$\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\sum_{o\in\mathcal{S}_{O/Y}^{l}}(h_{o}^{l-1}-m_{o}^{l})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{o}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}_{O/Y}^{l}|}.$
@@ A.2 Training of RGNN
$\bar{d}_{Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{Y^{t+}}|}\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}\hat{h}_{j}^{L,t},\quad\bar{d}_{O/Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{O/Y^{t+}}|}\sum_{o\in\mathcal{S}^{t}_{O/Y^{t+}}}\hat{h}_{o}^{L,t},$ (12)
@@
$\ell$ $=\frac{1}{T}\sum_{t\in[T]}\ell^{t}=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (13)
The gradient of loss $\ell^{t}$ at step $t$ with respect to $\alpha$ is computed by:
$\frac{\partial\ell^{t}}{\partial\alpha^{l}}=\left\{\begin{aligned} 0,\quad\ell<=0\\ \sum_{\tau=0}^{t}\beta(1-\beta)^{t-\tau}f_{\alpha^{l}}(L,\mathcal{S}^{t}_{Y^{t+}}),\quad\ell>0\end{aligned}\right.,$ (14)
@@
$f_{\alpha^{l}}(L,\mathcal{S}^{t}_{Y^{t+}})$ $=\frac{\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}(h_{j}^{l-1,t}-m_{j}^{l,t})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{j}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}^{t}_{Y^{t+}}|}$ (15)
$\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\sum_{o\in\mathcal{S}_{O/Y}^{l}}(h_{o}^{l-1,t}-m_{o}^{l,t})\prod_{\iota=l+1}^{L}\mathbb{I}(v_{o}\in\mathcal{S}_{K}^{\iota})\alpha^{\iota}}{|\mathcal{S}_{O/Y}^{l}|}.$
The gradient of loss $\ell^{t}$ at step $t$ with respect to $\beta$ is computed by:
$\frac{\partial\ell^{t}}{\partial\beta}=\left\{\begin{aligned} 0,\quad\ell<=0\\ f_{\beta}(t,L,\mathcal{S}^{t}_{Y^{t+}}),\quad\ell>0\end{aligned}\right.,$ (16)
where
$f_{\beta}(t,L,\mathcal{S}^{t}_{Y^{t+}})$ $=\frac{\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}\sum_{\tau=1}^{t}(h_{j}^{L,\tau}-\hat{h}_{j}^{L,\tau-1})(1-\beta)^{t-\tau}}{|\mathcal{S}^{t}_{Y^{t+}}|}$ (17)
$\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\sum_{o\in\mathcal{S}^{t}_{O/Y^{t+}}}\sum_{\tau=1}^{t}(h_{o}^{L,\tau}-\hat{h}_{o}^{L,\tau-1})(1-\beta)^{t-\tau}}{|\mathcal{S}_{O/Y^{t+}}^{t}|}.$
@@ Appendix B Additional Experimental Details
Prompt for extracting keywords Instruction: Extract the entities exist in this text and then provide the wikipedia links for the entities exist in this text. Output the entities and their wikipedia links in the list format, e.g., [[’entity1’, ’link1’], [’entity2’, ’link2’]].
<Passage>
@@
Implemental details of training. We initialize $\alpha^{1}=1.0$ and $\beta=1.0$, and use gradient descent to update the GNN and RGNN. For each iteration, we compute the average gradient of all the training samples and update the parameters with a learning rate of 1.0. The training will stop when the absolute value of gradient is smaller than 0.001 or the loss keeps increasing for five consecutive iterations. As there is often the whole label set of supporting passages for the question but not the individual index for each subquestion, we select the one with the lowest semantic distance as the label for each subquestion and then remove this index label from the label set for subsequent subquestions. Since training the RGNN requires ground-truth subquestions, we manually generate subquestions for 5 questions sampled from the preset training data. The training samples of MuSiQue, IIRC, and 2WikiMQA datasets for the RGNN are displayed in Appendix G. We also display the instructions and prompt templates of all the methods in Appendix F.
Appendix C Qualitative Results
@@
GNN-Ret improves the retrieval coverage of supporting passages. We compare the retrieval process between Dense and GNN-Ret and display the results on Table 3. GNN-Ret can retrieve all of the supporting passages while Dense fails in both of these two cases. For the first question, Dense can retrieve the knowledge that ‘the director of file Hotel By The Hour is Rolf Olsen.’. Since it considers passages in isolation during retrieval, the supporting passage for the inquiry information (‘which country is Rolf Olsen from’) cannot be retrieved with a poor semantic distance. In contrast, GNN-Ret is able to retrieve both of the supporting passages about ‘the director of file Hotel By The Hour is Rolf Olsen’ and ‘Rolf Olsen was an Austrian actor’. Consequently, it can output the correct final answer. This is attributed to the fact that GNN-Ret takes relatedness between passages into account and thus allows the supporting passages for inquiry information to accept the effect of semantic distances from those for background information, thereby improving the retrieval coverage of supporting passages.
Question: Which country the director of film Hotel By The Hour is from? Ground-truth answer: Austria Dense Retrieved passages: (1) Hotel by the Hour (German title:) is a 1970 West German crime film directed by Rolf Olsen and starring Curd Jürgens, Andrea Rau and Corny Collins. (2) … Final answer: Rolf Olsen is from Germany. GNN-Ret Retrieved passages: (1) Hotel by the Hour (German title:) is a 1970 West German crime film directed by Rolf Olsen and starring Curd Jürgens, Andrea Rau and Corny Collins. (2) Rolf Olsen (26 December 1919 – 3 April 1998) was an Austrian actor. Final answer: Rolf Olsen is from Austria. Question: What is the date of death of the performer of song Goodbye Pork Pie Hat? Ground-truth answer: January 5, 1979 Dense Retrieved passages: (1) ‘Goodbye Pork Pie Hat’ is a jazz standard composed by Charles Mingus, originally recorded by his sextet in 1959 as listed below, and released on his album. (2) … Final Answer: February 17, 2017 GNN-Ret Retrieved passages: (1)‘Goodbye Pork Pie Hat’ is a jazz standard composed by Charles Mingus, originally recorded by his sextet in 1959 as listed below, and released on his album. (2) Charles Mingus Jr. (April 22, 1922 – January 5, 1979) was an American jazz double bassist, pianist, composer and bandleader. Final answer: January 5, 1979
@@
RGNN-Ret enhances the retrieval process for subquestions. For the second question ‘What other notable work did the creator of Shrek make?’ in Table 4, both SelfAsk and RGNN-Ret are able to correctly answer the first subquestion and generate the second-step subquestion ‘What other notable works did William Steig create?’. However, SelfAsk cannot retrieve the knowledge about ‘other work of William Steig’, which locates in another passage about the book ‘Doctor De Soto’. Using dense retrieval fails to retrieve this passage for QA. In contrast, RGNN-Ret retrieve this passage since it enhances the semantic distance by integrating with the small semantic distances from the previous step. Consequently, RGNN-Ret outputs the correct answer. This qualitative example demonstrates that the RGNN can enhance the retrieval over steps for subquestions and thus improve the accuracy for QA.
Appendix D Limitations and Discussions
Appendix E Ethics Statement
We hold a firm conviction that our proposed GNN-Ret and RGNN-Ret have the potential to yield substantial positive outcomes in real-world scenarios associated with retrieval-augmented generation (RAG). GNN-Ret and RGNN-Ret significantly enhanced the retrieval coverage of supporting passages, thereby significantly improving the explainability and reliability of LLMs. Given the widespread application of LLMs, we anticipate that the implementation of our proposed retrieval techniques will extend their benefits to an increasing number of individuals, thereby catering to a broader range of needs.
Contrasting the numerous potential benefits of utilizing LLMs, it is crucial to acknowledge the existence of certain negative impacts, such as the generation of hallucinatory outputs and the associated costs. Addressing the issue of hallucinations in LLMs requires ongoing efforts to improve retrieval accuracy and provide comprehensive supporting passages to LLMs. By ensuring that LLMs have access to a wider range of relevant information, the occurrence of hallucinatory responses can be mitigated. Additionally, cost reduction strategies can be implemented by substituting specific steps of GNN-Ret and RGNN-Ret, e.g., keyword extraction. The pretrained small LMs could be an effective way to reduce the costs for applications.
Appendix F Prompts for Experiments

### v2

@@ 5 Conclusion
In this paper, we proposed GNN-Ret, an effective method to enhance retrieval for QA of LLMs by exploiting the inherent relatedness between passages on a graph of passages. By extending GNN-Ret to multi-hop reasoning questions, we proposed RGNN-Ret, which enhances the retrieval for subquestions through the interconnection between graphs of passages across steps. The experiments clearly demonstrated the superiority of both GNN-Ret and RGNN-Ret over baselines, highlighting the effectiveness of leveraging relatedness between passages to enhance retrieval. From these advantages, we believe that the incorporation of graph-based representations and the exploitation of passage relatedness can open up new avenues of research in the field of LLMs in understanding structural documents and answering complex questions.
6 Limitations
Costs of message passing. We first analyze the statistics of the graph of passages in experiments and detail results in Table 5. In larger datasets, each node typically exhibits a higher number of neighbors, attributed to an increased likelihood of keyword sharing among more nodes. We calculate the density of the graph of passages, denoted as $D=\frac{2|\mathcal{E}|}{|\mathcal{V}|(|\mathcal{V}|-1)}$, for each dataset in our experiments, based on the established definition Pieterse and Black (2005), where $\mathcal{E}$ and $\mathcal{V}$ represent the edges and vertices of the graph, respectively. The densities of all graphs are significantly less than 1, underscoring the substantial sparsity observed in these graphs across our experimental datasets.
We adopt relevant nodes sampling (Section 3.2) to allow the relevant node set with only top $K$ smallest semantic distances to pass the information to their neighbors during message passing. Assuming an average number of neighbors $\mathcal{V}_{K}$ among these $K$ relevant nodes, the time complexity for message passing can be derived as $O(K|\mathcal{V}_{K}|)$, considering a single-layer GNN. With $K$ set to 5 or 10 in our experiments, and given the small empirical average number of edges per node $\frac{|\mathcal{E}|}{|\mathcal{V}|}$ reported in Table 5, the time complexity of message passing is neglectable compared with LLM queries.
Datasets Number of nodes ($|\mathcal{V}|$) Average number of edges ($\frac{|\mathcal{E}|}{|\mathcal{V}|}$) Density
2WikiMQA 9815 91.38 0.018
IIRC 21866 259.22 0.024
MuSiQue 20071 335.77 0.033
Quality 1104 26.00 0.047
Table 5: Statistics of graph of passages in experiments.
@@ A.1 Training of GNN
$\bar{d}_{Y}^{L}=\frac{1}{|\mathcal{S}_{Y}|}\sum_{j\in\mathcal{S}_{Y}}h_{j}^{L},$ (6)
$\bar{d}_{O/Y}^{L}=\frac{1}{|\mathcal{S}_{O/Y}|}\sum_{o\in\mathcal{S}_{O/Y}}h_{o}^{L},$ (7)
@@
We update the parameters of GNN using gradient descent, and the backward process is presented as below. With only partial nodes accepting messages from neighbors, we first define the received node set at each layer $l$ as follows:
$\mathcal{S}_{R}^{l}=\left\{\begin{aligned} \mathcal{S}_{O}\cap\mathcal{S}_{K}^{L},\quad l=L\\ \mathcal{S}_{K}^{l},\quad\text{else}\end{aligned}\right..$ (9)
For each node $i\in\mathcal{S}_{R}^{L}$ at the last layer $L$, the gradient of $\ell$ w.r.t. $h_{i}^{L}$ is:
$\frac{\partial\ell}{\partial h^{L}_{i}}=\left\{\begin{aligned} 1.0,\quad i\in\mathcal{S}_{Y}\\ -1.0,\quad\text{else}\end{aligned}\right..$ (10)
For the layer $l<L$, the gradient of $l$ w.r.t. $\alpha^{l}$ is presented as follows:
$\frac{\partial\ell}{\partial\alpha^{l}}=\sum_{j\in\mathcal{S}_{R}^{l}}\frac{\partial\ell}{\partial h^{l}_{i}}(h_{i}^{l-1}-m_{i}^{l}).$ (11)
Assuming that node $i\in\mathcal{T}_{k}^{l-1}$ at layer $l-1$ transfers the message to the node set $\mathcal{S}_{K,i}^{l}$ at layer $l$ (one node can transfers messages to multiple neighbors), we update the gradient of $\alpha^{l}$ w.r.t. $h^{l-1}_{i}$ at the previous layer $l-1$ as follows:
$\frac{\partial\ell}{\partial h^{l-1}_{i}}=\sum_{j\in\mathcal{S}_{K,i}^{l}}-\alpha^{l}\frac{\partial\ell}{\partial h_{j}^{l}}+\mathbb{1}(i\in\mathcal{S}_{K,i}^{l})\alpha^{l}\frac{\partial\ell}{\partial h_{j}^{l}}.$ (12)
By propagating the gradients backward to the first layer, we can obtain the gradients for each layer of $\alpha$.
@@ A.2 Training of RGNN
$\bar{d}_{Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{Y^{t+}}|}\sum_{j\in\mathcal{S}^{t}_{Y^{t+}}}\hat{h}_{j}^{L,t},$ (13)
$\bar{d}_{O/Y^{t+}}^{L,t}=\frac{1}{|\mathcal{S}^{t}_{O/Y^{t+}}|}\sum_{o\in\mathcal{S}^{t}_{O/Y^{t+}}}\hat{h}_{o}^{L,t},$ (14)
@@
$\ell$ $=\frac{1}{T}\sum_{t\in[T]}\ell^{t}$
$=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (15)
We update the parameters of RGNN using gradient descent, and the backward process is presented as below. With only partial nodes accepting messages from neighbors, we first define the recieved node set $\mathcal{S}_{R}^{l,t}$ at layer $l$ at step $t$ as follows:
$\mathcal{S}_{R}^{l,t}=\left\{\begin{aligned} \mathcal{S}_{O}^{t}\cap\mathcal{S}_{K}^{L,t},\quad l=L\\ \mathcal{S}_{K}^{l,t},\quad\text{else}\end{aligned}\right..$ (16)
For each node $i\in\mathcal{S}_{R}^{L,t}$ at the last layer $L$, the gradient of $\ell^{t}$ w.r.t. $\hat{h}_{i}^{L,t}$ is:
$\frac{\partial\ell^{t}}{\partial\hat{h}^{L,t}_{i}}=\left\{\begin{aligned} 1.0,\quad i\in\mathcal{S}_{Y^{t}}^{t}\\ -1.0,\quad\text{else}\end{aligned}\right..$ (17)
Therefore, the gradient of $\ell$ w.r.t. $\hat{h}_{i}^{L,t}$ is:
$\frac{\partial\ell}{\partial\hat{h}^{L,t}_{i}}=\sum_{\tau=t}^{T}\frac{\partial\ell^{\tau}}{\partial\hat{h}^{L,t}_{i}}=\frac{\partial\ell^{t}}{\partial\hat{h}^{L,t}_{i}}+\frac{\partial\ell}{\partial\hat{h}^{L,t+1}_{i}}(1-\beta),$ (18)
@@
$\frac{\partial\ell}{\partial\hat{h}^{L,T}_{i}}=\frac{\partial\ell^{T}}{\partial\hat{h}^{L,T}_{i}}.$ (19)
Each node $j\in\mathcal{V}$ at step $t$ transfers the message to the same node at the next step through $\beta$. Therefore, the gradient of loss $\ell$ with respect to $\beta$ is computed by:
$\frac{\partial\ell}{\partial\beta}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{|\mathcal{V}|}\sum_{j\in\mathcal{V}}\frac{\partial\ell}{\partial\hat{h}^{L,t}_{j}}(h_{j}^{L,t}-\hat{h}_{j}^{L,t-1}).$ (20)
Recall that $\hat{h}^{L,t}_{i}=\beta{h}^{L,t}_{i}+(1-\beta){h}^{L,t-1}_{i}$ for each node $i\in\mathcal{V}$. The gradient of loss $\ell$ with respect to ${h}^{L,t}_{i}$ is computed by:
$\frac{\partial\ell}{\partial{h}^{L,t}_{i}}=\beta\frac{\partial\ell}{\partial\hat{h}^{L,t}_{i}}.$ (21)
With $\frac{\partial\ell}{\partial{h}^{L,t}_{i}}$, we propagate the gradient backward to the GNN and update the parameter $\alpha^{l}$ layer by layer using (11) and (12).
@@ Appendix B Additional Experimental Details
Datasets processing. We measure all the methods on four different QA datasets, including multi-hop Wikipedia reasoning datasets: (1) MuSiQue Trivedi et al. (2022), (2) IIRC Ferguson et al. (2020), (3) 2WikiMQA Ho et al. (2020), and a single-hop multi-choice QA dataset: (4) Quality Pang et al. (2021). For each multi-hop dataset, we randomly sample 500 questions from the development set, which provides the ground-truth supporting passages. We use 20 of them to train the GNN and select the rest as test data. We use all the Wikipedia documents of these 500 questions as the retrieval corpus, which is significantly larger than that provided by datasets [^4] Prior works often use 10–20 paragraphs or documents as the retrieval corpus for each question Shao et al. (2023); Wang et al. (2024).. For Quality, we randomly sample 30 articles with overall corresponding 561 questions and use these articles as the retrieval corpus. For each document or article, we split it into multiple passages with a maximum token length of 200 Sarthi et al. (2024).
@@
Prompt for extracting keywords Instruction: Extract the entities exist in this text and then provide the wikipedia links for the entities exist in this text. Output the entities and their wikipedia links in the list format, e.g., [[’entity1’, ’link1’], [’entity2’, ’link2’]]. <Passage>
@@
Implemental details of training. We initialize $\alpha=0.1$ for GNN-Ret and $\alpha^{1}=1.0$ and $\beta=1.0$ for RGNN-Ret, respectively. We then use gradient descent to update the GNN and RGNN. For each iteration, we compute the average gradient of all the training samples and update the parameters with a learning rate of 1.0. The training will stop when the absolute value of gradient is smaller than 0.001 or the loss keeps increasing for five consecutive iterations. As there is often the whole label set of supporting passages for the question but not the individual index for each subquestion, we select the one with the lowest semantic distance as the label for each subquestion and then remove this index label from the label set for subsequent subquestions. Since training the RGNN requires ground-truth subquestions, we manually generate subquestions for 5 questions sampled from the preset training data. The training samples of MuSiQue, IIRC, and 2WikiMQA datasets for the RGNN are displayed in Appendix J. We also display the instructions and prompt templates of all the methods in Appendix I.
Appendix C Selection of $K$ and $O$.
In order to effectively train the GNN and RGNN models, we conduct grid search for hyperparameters $K$ and $O$, which determine the number of relevant nodes to be sampled and the size of the competitive node set, respectively. To assess the impact of different values of $K$ and $O$, we evaluate the accuracy of GNN-Ret and RGNN-Ret on the MuSiQue and IIRC datasets with varying $K$ and $O$. The average accuracy across these two datasets is represented by the green points in Fig. 4. Results demonstrate that both GNN-Ret and RGNN-Ret consistently achieve stable accuracy performance across different settings of $K$ and $O$. Upon analyzing the results, we observe that GNN-Ret achieves the highest accuracy when $K=5$ and $O=25$, while RGNN-Ret performs best with $K=5$ and $O=10$. GNN-Ret adopts a larger value of $O$ since it considers the whole label set during training, while RGNN-Ret only considers the labels for the current step and subsequent steps. Therefore, we select these settings of $K$ and $O$ to train the GNN and RGNN models for experiments reported in Table 1.
(a) MuSiQue
(b) 2WikiMQA
Appendix D Statistics of retrieval accuracy.
We explore the EM between the retrieved passages and the ground-truth passages and collect the exact-match number of test samples with varying numbers of supporting passages required in questions on MuSiQue and 2WikiMQA datasets. We do not evaluate the statistics of retrieval accuracy for RGNN-Ret without the specific ground-truth supporting passages for each subquestion. Results in Fig. 5 show that our proposed GNN-Ret achieves higher retrieval accuracy compared with SBERT on MuSiQue and 2WikiMQA datasets. For instance, GNN-Ret outperforms SBERT by 29 and 43 exact-match test samples for the questions that require 2 supporting passages. This demonstrates that the GNN indeed improves retrieval coverage of supporting passages and the retrieval prioritization of supporting passages for subsequent reasoning.
Token number for retrieval 1k 3k 5k 10k 20k
SBERT 18.3 23.8 22.5 22.5 27.9
GNN-Ret 20.4 25.8 25.8 26.7 31.9
Table 6: Accuracy of GNN-Ret and SBERT with varying token numbers for retrieval using Qwen/Qwen2-7B-Instruct on MuSiQue dataset.
Appendix E Ablation Studies of Token Number for Retrieval
We conduct experiments with varying numbers of tokens for retrieval on SBERT and GNN-Ret with the long-context model Qwen/Qwen2-7B-Instruct Yang et al. (2024), which can handle more than 100k tokens. Results in Table 6 show that our proposed GNN-Ret consistently outperforms SBERT with varying numbers of tokens for retrieval, demonstrating the robustness of our method for long contexts.
GNN-Ret Learned values of $\alpha$ MuSiQue Learned values of $\alpha$ 2WikiMQA
F1 EM Acc F1 EM Acc
1 layer $\alpha=[0.326]$ 31.1 17.5 27.3 $\alpha=[0.277]$ 48.1 33.5 46.9
2 layers $\alpha=[0.564,0.821]$ 31.5 16.3 27.3 $\alpha=[0.582,0.578]$ 47.0 33.3 46.5
3 layers $\alpha=[0.506,0.511,0.944]$ 30.8 16.3 26.7 $\alpha=[0.504,0.597,0.802]$ 48.6 34.4 47.5
Table 7: F1, EM, and Accuracy of GNN-Ret with varying layers on MuSiQue and 2WikiMQA datasets.
Appendix F Extending GNN-Ret to Multiple Layers
Our proposed GNNs are designed to be extensible across multiple layers. To explore this capability, we conducted additional experiments by increasing the number of GNN layers. The results, along with the final learned values of $\alpha$, are presented in Table 7. The results indicate that GNN-Ret maintains consistent performance across different layer configurations. This observation could be attributed to the fact that many questions in datasets require only two-hop supporting passages, suggesting that a single-layer GNN is adequately equipped to address these queries.
Method MuSiQue 2WikiMQA
IRCoT 27.5 44.6
IRCoT + RGNN-Ret 30.4 (+2.9) 47.5 (+2.9)
ITER-RETGEN 27.1 43.8
ITER-RETGEN + RGNN-Ret 29.8 (+2.7) 47.5 (+3.7)
Table 8: Accuracy of IRCoT (w. RGNN-Ret) and ITER-RETGEN (w. RGNN-Ret) on MuSiQue and 2WikiMQA datasets.
Appendix G Effectiveness of RGNN-Ret for other multi-hop answering methods
To further demonstrate the effectiveness of RGNN-Ret, we supplement experiments by employing it with other multi-hop baselines IRCoT and ITER-RETGEN. These methods utilize the generated reasons for retrieval, followed by the generation of the next-step reason or the final answer. Results in Table 8 show that RGNN-Ret consistently and significantly improves the accuracy performance for other multi-hop baselines, which demonstrates the effectiveness and adaptability of RGNN-Ret in enhancing the retrieval performance for multi-hop answering processes.
Appendix H Qualitative Results
@@
GNN-Ret improves the retrieval coverage of supporting passages. We compare the retrieval process between SBERT and GNN-Ret and display the results on Table 9. GNN-Ret can retrieve all of the supporting passages while SBERT fails in both of these two cases. For the first question, SBERT can retrieve the knowledge that ‘the director of file Hotel By The Hour is Rolf Olsen.’. Since it considers passages in isolation during retrieval, the supporting passage for the inquiry information (‘which country is Rolf Olsen from’) cannot be retrieved with a poor semantic distance. In contrast, GNN-Ret is able to retrieve both of the supporting passages about ‘the director of file Hotel By The Hour is Rolf Olsen’ and ‘Rolf Olsen was an Austrian actor’. Consequently, it can output the correct final answer. This is attributed to the fact that GNN-Ret takes relatedness between passages into account and thus allows the supporting passages for inquiry information to accept the effect of semantic distances from those for background information, thereby improving the retrieval coverage of supporting passages.
Question: Which country the director of film Hotel By The Hour is from? Ground-truth answer: Austria SBERT Retrieved passages: (1) Hotel by the Hour (German title:) is a 1970 West German crime film directed by Rolf Olsen and starring Curd Jürgens, Andrea Rau and Corny Collins. (2) … Final answer: Rolf Olsen is from Germany. GNN-Ret Retrieved passages: (1) Hotel by the Hour (German title:) is a 1970 West German crime film directed by Rolf Olsen and starring Curd Jürgens, Andrea Rau and Corny Collins. (2) Rolf Olsen (26 December 1919 – 3 April 1998) was an Austrian actor. Final answer: Rolf Olsen is from Austria. Question: What is the date of death of the performer of song Goodbye Pork Pie Hat? Ground-truth answer: January 5, 1979 SBERT Retrieved passages: (1) ‘Goodbye Pork Pie Hat’ is a jazz standard composed by Charles Mingus, originally recorded by his sextet in 1959 as listed below, and released on his album. (2) … Final Answer: February 17, 2017 GNN-Ret Retrieved passages: (1)‘Goodbye Pork Pie Hat’ is a jazz standard composed by Charles Mingus, originally recorded by his sextet in 1959 as listed below, and released on his album. (2) Charles Mingus Jr. (April 22, 1922 – January 5, 1979) was an American jazz double bassist, pianist, composer and bandleader. Final answer: January 5, 1979
@@
RGNN-Ret enhances the retrieval process for subquestions. For the second question ‘What other notable work did the creator of Shrek make?’ in Table 10, both SelfAsk and RGNN-Ret are able to correctly answer the first subquestion and generate the second-step subquestion ‘What other notable works did William Steig create?’. However, SelfAsk cannot retrieve the knowledge about ‘other work of William Steig’, which locates in another passage about the book ‘Doctor De Soto’. Using SBERT for retrieval fails to retrieve this passage for QA. In contrast, RGNN-Ret retrieve this passage since it enhances the semantic distance by integrating with the small semantic distances from the previous step. Consequently, RGNN-Ret outputs the correct answer. This qualitative example demonstrates that the RGNN can enhance the retrieval over steps for subquestions and thus improve the accuracy for QA.
Appendix I Prompts for Experiments